OpenStudy (anonymous):

How do you find integrating factor of differential equation ? Help please!!!

OpenStudy (anonymous):

This is the equation dy/dt = −2y + 2t − 1

OpenStudy (unklerhaukus):

integrating factor for firs order DE's \[y'+p(x)y=q(q) \]is \[e^{\intp(x)dx)\] and you dont have to worry about the constant

OpenStudy (unklerhaukus):

ops \[e^{\int p(x)dx}\]

OpenStudy (unklerhaukus):

\[dy/dt = −2y + 2t − 1 \] \[y'+2y=2t-1\] so you p(x) is 2 integrating factor is \[=e^{\int2dx}\] \[=e^{2x}\]

OpenStudy (anonymous):

thank you soo much!!! if you're still there are you able to help me with one more!! :( i'm stuck!

OpenStudy (anonymous):

dy/dt− 2y = 3e^(2t)